The flywheel of an engine has moment of inertia 2.50kg.m^2 about its rotation axis.What constant torque is required to bring it up to an angular speed of 400rev/min in 8s,starting from rest
1
Expert's answer
2020-11-16T07:40:29-0500

According the the law of rotational movement, the torque is


"\tau = I\varepsilon"

where "I = 2.5\space kg\cdot m^2" and "\varepsilon" is the angular acceleration.

By definition, the angular acceleration is:

"\varepsilon = \dfrac{\omega - \omega_0}{t} = \dfrac{\omega}{t}"

where "\omega = 2\pi \cdot 400 = 800\pi \dfrac{rad}{min} = \dfrac{40}{3}\pi\space \dfrac{rad}{s}" is the final angular velocity and "\omega_0 = 0" is the intial angular velocity (since starts from rest). Here "t = 8s" is the time of velocity increasing.

Substituting angular acceleration, obtain:


"\tau = I\varepsilon = \dfrac{I\omega}{t}\\
\tau = \dfrac{2.5\cdot \dfrac{40}{3}\pi}{8}\approx 13.1\space N\cdot m"

Answer. "13.1\space N\cdot m".